41. A revenue expenditure A) B) Increases the book value of a long-term asset Decreases the book value of a long-term asset Increases a revenue account Increases an expense account D) 42. Which of the following is a capital expenditure? A) B) Cost to add air conditioning to a company car Cost to purchase garbage cans for the company conference room Cost to replace spark plug on company lawnmower D) Cost to have store windows washed 43. Which of the following is a revenue expenditure? A) B) C) D) Adding air conditioning to a company car Adding a loading and unloading dock to the companys warehouse Purchasing a microcomputer for the marketing department Replacing the battery in a company truclk45. At January 1, 2016, Egret Company had total assets of S1,800,000 and at December 31, 2016, its total assets were $2,200,000. Egrets net sales for 2016 were $1,700,000 and its 2016 net income was $110,000. Egrets return on assets for 2016 is: A) B) C) D) 5.5% 5.0% 6.5% 85.0%
Answer #1
41
A revenue expenditure increases an expense account.
Option D is correct
42
Cost to add air conditioning to a company car is a Capital expenditure, as it is expected to benefit more than one accounting period
Option A is correct
43
Replacing the battery in a Computer truck is a revenue expenditure, as it is normal replacement cost.
Option D is correct
44
Average total assets = (1800000+2200000)/2 = $2000000
Return on assets = Net income/Average total assets = 110000/2000000= 5.5%
Option A is correct